OpenStudy (anonymous):

The sales totals at Robert's food store have increased linearly over the month. Which of these best shows the sales in the first three months?

OpenStudy (danjs):

Hi, welcome to open study!

OpenStudy (anonymous):

(A)$1200 in the first month, $1212 in the second month, $1224.12 in the third month (B)$1200 in the first month, $1236 in the second month, $1273.08 in the third month (C)$1200 in the first month, $1285.48 in the second month, $1370.96 in the third month (D)$1200 in the first month, $1224.28 in the second month, and $1248.48 in the third month

OpenStudy (anonymous):

lol i was here on openstudy for like many months

OpenStudy (anonymous):

I think the answer is C because there is a constant difference of $24.28

OpenStudy (danjs):

When the question says that the sales increased linearly, it means that sales will increase the same amount for each increase in time

OpenStudy (danjs):

sorry, said you were new here. haha

OpenStudy (danjs):

oh linear algebra. lol ok

OpenStudy (anonymous):

its alright lol

OpenStudy (danjs):

Right, the answer here would be the one that increased the same amount from month to month

OpenStudy (danjs):

the increase per month would be the slope of the line

OpenStudy (anonymous):

Thanks for clarifying :)

OpenStudy (danjs):

welcom.. anytime it is linear, think of a line and a slope, the increase per unit is the same... y = k*x ---y is some constant times the x value

OpenStudy (anonymous):

okay i'll remember and i know that equation too. i've seen that everywhere

OpenStudy (danjs):

yeah, y=k*x is the probs where they say y is directly proportional to x. the only difference between y = k*x and y=m*x + b, is y=k*x always goes through the origin, where as y=mx+b can intercept the y-axis anyehere at b.

OpenStudy (danjs):

any line is a y=k*x direct proportion, the direct proportion is the slope

OpenStudy (anonymous):

alright got it

OpenStudy (arabpride):

Please close this post if you're done. Thanks! :D